QUESTION 5 Demand and supply of baby food are given by QAD-10-2P and Qs-2P respectively. What is the equilibrium price in the market? $4/unit $6/unit $9/unit None of the above QUESTION6 In Question 5, what is the equilibrium quantity supplied in the market? 4 units 3 units 5 units None of the above
Answer #1

5. Qd= 10-2P

Qs= 2P

10-2P = 2P

4P= 10

P= $ 2.5/unit. (Equilibrium price)

Hence, option(D) i.e none of the above is correct.

6. Qs = 2P

Qs = 2(2.5) =5 units.   (Equilibrium quantity supplied)

Hence,option(C) is correct.